# Bayesian power spectral density estimation for LISA noise based on P-splines with a parametric boost

### v1

@@ III Simulation study
where $\sigma^{2}$ is the variance and $(a_{1},a_{2},a_{3},a_{4})$ are the AR model parameters. We set $f_{s}=1$, the variance to unity, and $a_{1}=0.9$, $a_{2}=-0.9$, $a_{3}=0.9$, $a_{4}=-0.9$, following previous work [18, 19]. These parameters produce a spectrum with two sharp peaks (Figure 1(a)). Estimation accuracy is measured using the integrated absolute error,
@@
We simulate series of length $n\in\{128,256,512\}$, generating $500$ independent realizations for each $n$. For each realization, we estimate the log-spectral density by running $20,000$ iterations with a $1,000$-iteration burn-in, and a thinning factor of $10$, resulting in $1,900$ samples for posterior inference. We place $20$ knots using the quantile-based method here, applied to the periodogram and parametric PSD ratio. We calculate the covariance matrix from a large posterior sample to initialize the AMH algorithm across the simulations to generate the posterior samples.
@@
Figure 1: Comparison of PSD estimation (a) and IAE distributions (b) for Model 1 (green) and Model 2 (orange) parametric models.
@@
We simulated 150 datasets of length $n=512$ and estimated the PSD using Models 1 and 2 with different numbers of knots ($5$, $10$, and $20$). The corresponding IAEs were then computed, and the results are shown in Figure 2. The IAEs for Model 2 remain relatively stable across knot values, indicating little residual structure after the parametric fit. Whereas Model 1 requires a larger number of knots to match the accuracy of Model 2. These results indicate that a well-specified parametric model substantially reduces the number of knots needed for the P-spline correction to achieve comparable accuracy when no suitable parametric model is available.
Figure 2: IAE distribution with various numbers of knots for two semi-parametric models for $n=512$. Boxes indicate the interquartile range, whiskers the full range, and horizontal lines the median.
@@ IV Application to LISA Noise
We demonstrate the application of the blocked semi-parametric P-spline framework to second-generation TDI-$X$ channel data dominated by low-frequency test-mass acceleration (TM) and high-frequency optical metrology readout (OMS) noise. We analyze simulated LISA noise timeseries from the public Zenodo repository (DOI: 10.5281/zenodo.15698080), generated using LISA Instrument v1.1.1 [25, 26].
@@ IV.1 Noise Models
The simulation implements LDC Spritz noise curves under simplified orbital dynamics with fixed 8.3-second light travel times for inter-spacecraft laser beam propagation. Second-generation Michelson TDI combinations are computed via PyTDI [27, 28]. The simulated noise PSD $S_{\rm X}(f)$ can be approximated using analytical models for test-mass acceleration noise [26]
@@
where $a_{\rm TM}=2.4\times 10^{-15}\mathrm{ms}^{-2}$, $f_{1}=4\times 10^{-4}\,\mathrm{\mathrm{Hz}}$, $f_{2}=8\times 10^{-3}\,\mathrm{\mathrm{Hz}}$, and readout noise [26]
@@
where $a_{\rm OMS}=7.9\times 10^{-12}\mathrm{ms}^{-2}$, $f_{3}=2\times 10^{-3}\,\mathrm{Hz}$. The TM and OMS noise components are given in units of Hz (see the Zenodo data release for more details on the data generation [25]). Taking into account the transfer functions through TDI [29], the theoretical approximate PSD for the $X$ channel is given by
@@ IV.2 Application of P-spline framework
Figure 3: Power-spectral-density (PSD) estimation for the second-generation Michelson TDI $X$ channel using the OMS parametric noise model. Results are shown for data sets of 3 months, 6 months, and 1 year (light-to-dark red). Top: shaded regions show the middle $90\%$ credible interval (CI) for each observing time. The dashed black curve is the theoretical PSD $S_{\rm X}$, and the gray trace is a representative 5-day block periodogram. Bottom: relative error of the CI of log PSD against $\log(S_{\rm X})$. Longer observation times yield narrower CIs and smaller relative errors.
We partition LISA noise time series of duration $T\in\{3,6,12\}$ months into $J\in\{18,36,73\}$ five-day segments and apply a Kaiser window with $\beta=5$ (following noise-4a dataset settings [25]) to each segment. We then apply the blocked likelihood with 15 equidistant logarithmically-spaced frequency knots to the segmented periodogram.
@@
Figure 3 compares PSD estimates for three observing durations (3, 6, and 12 months). The top panel displays $90\%$ uniform credible intervals (CIs) for each duration, with darker orange shading corresponding to longer observing periods. The dashed black curve displays $S_{\rm X}(f)$, while the gray trace shows a representative 5-day block periodogram. As observing time increases, the CIs contract across the LISA band, reflecting the reduced uncertainty from additional data. The posterior median closely tracks $S_{\rm X}(f)$ despite the misspecified parametric component, demonstrating that the P-splines effectively correct for the parametric model inadequacies. Moreover, the analysis yields precise estimates in the central millihertz region where LISA exhibits maximum sensitivity to a stochastic gravitational wave background [30, 31].
@@
The approach consistently yields minimal uncertainties in the central millihertz band—precisely where stochastic gravitational wave background detection sensitivity is maximized [32]. This performance remains robust across all timescales tested, supporting the framework’s applicability for LISA noise characterization across mission-relevant timescales.
@@ V Discussion
We have developed a Bayesian semiparametric method for estimating the PSD of a stationary time series, focusing on LISA noise characterization. The PSD is expressed as the geometric mean of a parametric model and a nonparametric P-spline correction, combining the efficiency of a well-specified parametric form with the flexibility and robustness of a nonparametric approach. A logarithmically spaced frequency grid ensures resolution where the data are most informative. At the same time, quantile-based knot placement focuses model complexity on regions with the most significant deviations from the parametric PSD. A hierarchical penalty prior regularizes the spline coefficients, mitigating overfitting without compromising adaptability.
@@
The method has several practical advantages for LISA analysis. It can accommodate partial parametric knowledge without requiring the full PSD to be specified a priori. It remains computationally feasible for year-long data sets, enabling repeated use in iterative global-fit analysis pipelines [33, 34, 35]. Moreover, its nonparametric correction can capture subtle deviations from nominal instrument models, making it a valuable diagnostic tool.
There are also natural extensions. GPU acceleration should substantially reduce runtimes, making the method suitable for large-scale simulations and Monte Carlo studies. Extending the framework to separate stochastic gravitational wave backgrounds from instrument noise [36] would enable joint noise and signal inference. Generalising the spline correction to cross-spectral density matrices would allow simultaneous modelling of the correlated $X$, $Y$, and $Z$ channels, providing a more complete representation of the LISA data. Beyond LISA, the method could be adapted to estimate PSDs in ground-based detectors such as those in the LIGO-Virgo-KAGRA network, where it may assist in rapid noise modelling and continuous-wave searches. Using spline correction for cross-spectral density matrices could also be useful if correlated noise is present in the Einstein Telescope triangular configuration [37]. An additional avenue is developing a time–frequency formulation that enables the estimation of evolving PSDs in the presence of non-stationary noise, which is relevant for both LISA and terrestrial detectors.
@@
The software developed for this project is open-source and publicly available from the GitHub repository https://github.com/nz-gravity/npc.git, which contains all source code, example scripts, and configuration files needed to reproduce the results. Installation instructions and dependency information (tested on Python $\geq$3.10)) are provided in the repository README. The software is released under the MIT License, permitting free use, modification, and distribution. We also plan to release a PyPI package, fully compatible with JAX and GPU acceleration, allowing for easy installation and efficient execution on modern hardware. The datasets generated are available available at the public Zenodo repository (DOI: 10.5281/zenodo.17169792). LISA time series data is available at DOI: 10.5281/zenodo.15698080.
@@ Acknowledgements.
We thank Quentin Baghi, John-Baptiste Bayle, Matt Edwards, Martina Muratore, and others in the LISA Non-Stationarities Group for helpful discussions. We additionally thank John-Baptiste Bayle for providing the LISA dataset, and the University of Glasgow for the computing resources that supported JB in simulating the dataset. NA, NC, PMR, RM, and AV gratefully acknowledge support from the Marsden Fund Council grant MFP-UOA2131, funded by the New Zealand Government and managed by the Royal Society Te Apārangi. For this project, NC has received financial support from the CNRS through the MITI interdisciplinary programs, and the French Agence Nationale de la Recherche. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government.

### v2

@@ III Simulation study
where $\sigma^{2}$ is the variance and $(a_{1},a_{2},a_{3},a_{4})$ are the AR model parameters. We set $f_{s}=1$, the variance to unity, and $a_{1}=0.9$, $a_{2}=-0.9$, $a_{3}=0.9$, $a_{4}=-0.9$, following previous work [edwards_bayesian_2019, maturana-russel_bayesian_2021]. These parameters produce a spectrum with two sharp peaks (Figure 1(a)). Estimation accuracy is measured using the integrated absolute error,
@@
We simulate series of length $n\in\{128,256,512\}$, generating $500$ independent realizations for each $n$. For each realization, we estimate the log-spectral density by running $20,000$ iterations with a $1,000$-iteration burn-in, and a thinning factor of $10$, resulting in $1,900$ samples for posterior inference. We calculate the covariance matrix from a large posterior sample to initialize the AMH algorithm across the simulations to generate the posterior samples.
@@
For both models, we place $20$ knots using the quantile-based method applied to the ratio of the periodogram and the parametric PSD. Using the same number of knots ensures a fair comparison between the models, isolating differences due to the parametric component rather than spline resolution.
@@
(c) IAE distributions for different numbers of knots for $n=512$. Boxes indicate the interquartile range, whiskers the full range, and horizontal lines the median.
Figure 1: Performance comparison between semi-parametric PSD estimators. (a) PSD estimates for a single realization; (b) IAE distributions across dataset sizes; (c) Sensitivity of IAE to the number of knots. Model 1 (green) and Model 2 (orange) are shown throughout.
@@
To assess the impact of $N_{\xi}$ on estimation accuracy, we analyze 50 simulated AR(4) datasets of length $n=512$ and estimate the PSD using Models 1 and 2 with $N_{\xi}=\{5,10,15,20,25\}$. Figure 1(c) displays the IAE as a function of $N_{\xi}$ for both models. For Model 2, the IAE remains largely unchanged across the tested range, indicating that the hierarchical roughness-penalty prior effectively guards against overfitting even when the parametric component captures nearly all spectral structure. In contrast, Model 1 shows a reduction in IAE as $N_{\xi}$ increases from $5$ to approximately $10$, beyond which additional knots yield only marginal improvements. These results demonstrate that a well-specified parametric model reduces the number of knots needed for the P-spline correction to achieve a given accuracy. When no suitable parametric approximation is available, more knots are required, but the non-parametric correction still performs well.
@@ IV Application to LISA Noise
We demonstrate the application of the blocked semi-parametric P-spline framework to second-generation TDI-$X$ channel data dominated by low-frequency test-mass acceleration (TM) and high-frequency optical metrology readout (OMS) noise. We analyze simulated LISA noise timeseries from the public Zenodo repository (DOI: 10.5281/zenodo.15698080), generated using LISA Instrument v1.1.1 [bayle_lisa_2025, bayle_unified_2023].
@@ IV.1 Noise Models
The simulation implements LDC Spritz noise curves under simplified orbital dynamics with fixed 8.3-second light travel times for inter-spacecraft laser beam propagation. Second-generation Michelson TDI combinations are computed via PyTDI [staab_pytdi_2025, tinto_second-generation_2023]. The simulated noise PSD $S_{\rm X}(f)$ can be approximated using analytical models for test-mass acceleration noise [bayle_unified_2023]
@@
where $a_{\rm TM}=2.4\times 10^{-15}\mathrm{ms}^{-2}$, $f_{1}=4\times 10^{-4}\,\mathrm{\mathrm{Hz}}$, $f_{2}=8\times 10^{-3}\,\mathrm{\mathrm{Hz}}$, and readout noise [bayle_unified_2023]
@@
where $a_{\rm OMS}=7.9\times 10^{-12}\mathrm{ms}^{-2}$, $f_{3}=2\times 10^{-3}\,\mathrm{Hz}$. The TM and OMS noise components are given in units of Hz (see the Zenodo data release for more details on the data generation [bayle_lisa_2025]). Taking into account the transfer functions through TDI [quang_nam_time-delay_2023], the theoretical approximate PSD for the $X$ channel is given by
@@ IV.2 Application of P-spline framework
Figure 2: Power-spectral-density (PSD) estimation for the second-generation Michelson TDI $X$ channel using the OMS parametric noise model. Results are shown for data sets of 3 months, 6 months, and 1 year (light-to-dark red). Top: shaded regions show the middle $90\%$ credible interval (CI) for each observing time. The dashed black curve is the theoretical PSD $S_{\rm X}$, solid purple curve is the OMS parametric PSD, and the gray trace is a representative 5-day block periodogram. Bottom: relative error of the CI of log PSD against $\log(S_{\rm X})$. Longer observation times yield narrower CIs and smaller relative errors.
We partition LISA noise time series of duration $T\in\{3,6,12\}$ months into $J\in\{18,36,73\}$ five-day segments and apply a Kaiser window with $\beta=5$ (following noise-4a dataset settings [bayle_lisa_2025]) to each segment. We then apply the blocked likelihood with 15 equidistant logarithmically-spaced frequency knots to the segmented periodogram.
@@
Figure 2 compares PSD estimates for three observing durations (3, 6, and 12 months). The top panel displays $90\%$ uniform credible intervals (CIs) for each duration, with darker orange shading corresponding to longer observing periods. The dashed black curve displays $S_{\rm X}(f)$, while the gray trace shows a representative 5-day block periodogram. The purple curve shows our parametric model (OMS parametric PSD). As observing time increases, the CIs contract across the LISA band, reflecting the reduced uncertainty from additional data. The posterior median closely tracks $S_{\rm X}(f)$ despite the misspecified parametric component, demonstrating that the P-splines effectively correct for the parametric model inadequacies. Moreover, the analysis yields precise estimates in the central millihertz region where LISA exhibits maximum sensitivity to a stochastic gravitational wave background [Christensen_2018, babak_lisa_2021].
@@
The approach consistently yields minimal uncertainties in the central millihertz band—precisely where stochastic gravitational wave background detection sensitivity is maximized [robson_construction_2019]. This performance remains robust across all timescales tested, supporting the framework’s applicability for LISA noise characterization across mission-relevant timescales.
@@ V Discussion
We have developed a Bayesian semiparametric method for estimating the PSD of a stationary time series, focusing on LISA noise characterization. The PSD is expressed as the geometric mean of a parametric model and a nonparametric P-spline correction, combining the efficiency of a well-specified parametric form with the flexibility and robustness of a nonparametric approach. Non-uniform knot placement focuses model complexity on regions where the data are most informative. A hierarchical penalty prior regularizes the spline coefficients, mitigating overfitting without compromising adaptability.
@@
The method has several practical advantages for LISA analysis. It can accommodate partial parametric knowledge without requiring the full PSD to be specified a priori. It remains computationally feasible for year-long data sets, enabling repeated use in iterative global-fit analysis pipelines [Littenberg:2023xpl, Katz:2024oqg, Deng:2025wgk]. Moreover, its nonparametric correction can capture subtle deviations from nominal instrument models, making it a valuable diagnostic tool.
There are also natural extensions. GPU acceleration should substantially reduce runtimes, making the method suitable for large-scale simulations and Monte Carlo studies. Extending the framework to separate stochastic gravitational wave backgrounds from instrument noise [Boileau:2021sni] would enable joint noise and signal inference. Generalising the spline correction to cross-spectral density matrices would allow simultaneous modelling of the correlated $X$, $Y$, and $Z$ channels, providing a more complete representation of the LISA data. Beyond LISA, the method could be adapted to estimate PSDs in ground-based detectors such as those in the LIGO-Virgo-KAGRA network, where it may assist in rapid noise modelling and continuous-wave searches. Using spline correction for cross-spectral density matrices could also be useful if correlated noise is present in the Einstein Telescope triangular configuration [Janssens:2022xmo]. An additional avenue is developing a time–frequency formulation that enables the estimation of evolving PSDs in the presence of non-stationary noise, which is relevant for both LISA and terrestrial detectors.
@@
The software developed for this project is open-source and publicly available from the GitHub repository https://github.com/nz-gravity/npc.git, which contains all source code, example scripts, and configuration files needed to reproduce the results. Installation instructions and dependency information (tested on Python $\geq$3.10)) are provided in the repository README. The software is released under the MIT License, permitting free use, modification, and distribution. We also plan to release a PyPI package, fully compatible with JAX and GPU acceleration, allowing for easy installation and efficient execution on modern hardware. The datasets generated are available available at the public Zenodo repository [aimen_bayesian_2025]. LISA time series data is available at [bayle_lisa_2025].
@@ Acknowledgements.
We thank Quentin Baghi, Jean-Baptiste Bayle, Matt Edwards, Martina Muratore, and others in the LISA Noise Non-Stationarities Group (part of the “Deep analysis group” of the Distributed Data Processing Centre, DDPC) for helpful discussions. We additionally thank Jean-Baptiste Bayle for providing the LISA dataset, and the University of Glasgow for the computing resources that supported JB in simulating the dataset. NA, NC, PMR, RM, and AV gratefully acknowledge support from the Marsden Fund Council grant MFP-UOA2131, funded by the New Zealand Government and managed by the Royal Society Te Apārangi. For this project, NC has received financial support from the CNRS through the MITI interdisciplinary programs, and the French Agence Nationale de la Recherche. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government.